OpenStudy (peterg2001):

help please @Jhannybean All of the following are equivalent except _____. 5 is to 3 45/27 15/5 7.5:4.5

OpenStudy (jhannybean):

\[\frac{45}{27}\] if you divide both top and bottom by 9, what so you get?

OpenStudy (jhannybean):

\[7.5 = 7\frac{1}{2} = \frac{15}{2}\]\[4.5 = 4\frac{1}{2} = \frac{9}{2}\]\[\frac{\dfrac{15}{2}}{\dfrac{9}{2}} = \frac{15}{2}\cdot \frac{2}{9} = \frac{15}{9}\]Divide \(\dfrac{15}{9}\) by 3 both bottom and top, what do you get?

OpenStudy (jhannybean):

Then analyze \(\dfrac{15}{5}\). Can you find a number that will divide both top and bottom to give you \(\dfrac{5}{3}\)?

OpenStudy (jhannybean):

By putting each of your answer choices in fraction form and then reducing them, you should get \(\dfrac{5}{3}\). If you do not get \(\dfrac{5}{3}\) that will be your answer.